OpenStudy (anonymous):

the cost of $60,000 cost of a certain charter flight must be shared equally by those who take the trip. a certain number of travelers have already signed up for the trip. if 100 additional customers sign up, the cost toe of each of the original patrons will be reduced by $100. how many have already signed up?

OpenStudy (anonymous):

Let t be the number of travelers who have already signed up for the trip. Let c be the cost for the trip for each t. From the problem statement, the following equations can be derived.\[\left\{\frac{60000}{t}=c,\frac{60000}{t+100}=c-100\right\}\]@nushin Can you solve the equtions for t, the number of traelers who hve already signed up?

OpenStudy (anonymous):

no. can you please show me how its done step by step?

OpenStudy (anonymous):

Refer to the first equation. c = 60000/t . Replace c in the second equation with 60000/t . You now have one equation in one unknown.\[\frac{60000}{t+100}\text{=}\frac{60000}{t}-100 \]Can you proceed from here and find t ?

OpenStudy (anonymous):

yes i can. Thank you so much. :)

OpenStudy (anonymous):

Thank you for the medal.